OpenStudy (anonymous):

anybody in algebra 1 and is taking solving equations unit test ?

OpenStudy (anonymous):

I used to be in algebra 1

OpenStudy (anonymous):

you wanna help me with this test ?

OpenStudy (anonymous):

Sure. Hopefully I can answer all your questions :)

OpenStudy (anonymous):

okay so first question (2h-8)-h=h-16 h=8 h=-8 infinitely many solutions no soulutions

OpenStudy (anonymous):

(2h-8)-h=h-16 add h to bother sides 2h-8 = 2h-16 add 8 to both sides 2h = 2h -8 and so there are no solutions. questions?

OpenStudy (anonymous):

*both not bother

OpenStudy (anonymous):

6(z-2)+4=10 z=2 z=3 z=-3 z=-4

OpenStudy (ttp):

@vonnaj are you in FLVS?

OpenStudy (anonymous):

solve using distributive property 6z-12+4=10 combine like terms 6z-8=10 do the opposite and add ten on both sides 6z=18 divide both sides by 6 z=3